OpenStudy (anonymous):

Show all steps used to solve the problem. One point for steps and one point for solution. Be sure that you check your solution for correctness. If 3a - 9 = 6, what is the value of 4a - 1?

OpenStudy (istim):

You first must solve for 3a-9=6.

OpenStudy (istim):

In which you are looking for "a". Can you do that for me? It's algebra.

OpenStudy (istim):

Hello?

OpenStudy (anonymous):

i know its algebra thats what im doing

OpenStudy (istim):

Ok. That's good. Did you solve it yet then?

OpenStudy (anonymous):

no

OpenStudy (istim):

Ok. Well. First you need to isolate a.

OpenStudy (istim):

And we do the reverse of PEDMAS or BEDMAS to do so.

OpenStudy (istim):

That means we do the addition and subtraction before multiplication and division.

OpenStudy (istim):

You with me?

OpenStudy (anonymous):

yes

OpenStudy (istim):

Ok. You know a little bit of what you're suppose to do?

OpenStudy (anonymous):

on this question no

OpenStudy (istim):

Ok. Well, let's add 9 to both sides. Can you do that for me?

OpenStudy (istim):

Hey, you still there?

OpenStudy (anonymous):

this is concuseing cause i have no paper

OpenStudy (istim):

You should get some. That would make this process much easier.

OpenStudy (istim):

Um, what's concuseing?

OpenStudy (anonymous):

no really?? If i had some i would

OpenStudy (istim):

No scrap paper? Oh, you're doing an online course?

OpenStudy (istim):

You can do it directly here if you'd like, with the draw or equation button below your communication text box.

OpenStudy (anonymous):

3a - 9 = 6 value of 4a - 1 3a − 9 = 6 3a = 6 + 9 3a = 15 a = 15/3 a = 5 4a − 1 = 4(5) − 1 = 19

OpenStudy (istim):

Looks good to me. Nice work.

OpenStudy (anonymous):

thanks

OpenStudy (istim):

Seems like you understand your subject after all.

OpenStudy (anonymous):

not really

OpenStudy (istim):

Well, to a certain degree at the very least.

OpenStudy (istim):

Nice work. Good luck with your other stuff.

OpenStudy (anonymous):

thankz